Suppose you buy a round lot of Francesca Industries stock (100 shares) on 65 percent margin when the stock is selling at $15 a share. The broker charges a 8 percent annual interest rate, and commissions are 4 percent of the stock value on the purchase and sale. A year later you receive a $0.75 per share dividend and sell the stock for $21 a share. What is your rate of return on Francesca Industries? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Solution :-

Stock Value = 100 Shares * $ 15 = $1500

Initial Margin = $1500 * 65% = $975

Commission on Purchase = $1500 * 4% = $60

Interest Charged For one year = 1500 * 35% * 8% = $42

Commision on Sales = 100 Shares * $21 * 4% = $84

Dividend Received = 100 Shares * $0.75 = $75

Increment in Price Profit = 100 Shares * ($21 - $15) = $600

Net Income on Sale = Dividend + Increment in Price - Commission on Sale - Interest = $75 + $600 - $120 - $42 = $513

Net Investment = Initial Margin + Commision on Purchase = $975 + $60 = $1035

Now Rate of Return = $513 / $1035 = 47.72%